Thirty-three kilograms of nickel that was 35% pire were mixed with a nickel alloy that was 21% pure. How many kilograms of each kind were used to produce an alloy that was 27% pure?
Answer by venugopalramana(3286)

Thirty-three kilograms of nickel that was 35% pire were mixed with a nickel alloy that was 21% pure. How many kilograms of each kind were used to produce an alloy that was 27% pure?
LET 21% NICKEL USED = X KG
WE GET X+33 KG OF 27% NICKEL ALLOY
NICKEL IN 35% PURE ALLOY = 33*35/100
NICKEL IN 21% NICKEL ALLOY = X*21/100
NICKEL IN 27% NICKEL ALLOY = (X+30)*27/100
SO.....
33*35/100 + X*21/100 = (X+30)*27/100
1155+21X=27X+810
27X-21X=1155-810=345
6X=345
X=57.5 KG